What is the probability that your birthday woll fall on a Saturday or Sunday?
It depends on where in the world you are. In many countries maternity teams are less keen to work weekends. This is partly because there are fewer ancillary support teams, the maternity teams themselves want a home life and so on. So, to some extent, weekend births are rarer than would be expected purely by chance.
What is the probability of my house burning down?
The answer depends on very many variable which have not been specified. Most important is what your house is made of. Do you have a fire alarms and fire extinguishers? Are there any combustible materials stored in the house? Do you smoke? Do you deep-fry food a lot?
What do results tell you with regard to whether the data is normally distributed?
Your question may be a little overly complex. Usually you start with some degree of confidence that the variable(s) you are studying are normally distributed; you don't do an experiment and wait for the results to tell you. Your confidence that a variable is normally distributed helps you to determine whether or not your results are significantly different from what you would expect by chance. If I have missed the meaning of your question, my apologies.
What is the probability of rolling a 1 or 5 on a die?
you dont know that well it depends how many sides does it have lol if it is 8 it is likely if its 6 its mostlikly lol y do u care
What is the proper name given to the law that state If something can go wrong it will go wrong?
Murphy's Law.
What is the probability of drawing 2 cards that are all 6 from a deck of cards?
The probability of drawing 2 sixes from a deck of 52 cards is (4 in 52) times (3 in 51) which is (12 in 2652) or (1 in 221) or about 0.004525.
How many diamonds in a deck of cards are greater than 10?
Four; the Jack, Queen, King, and Ace of Diamonds.
What test provides a statement of probability that experimental outcome was due to random variation?
statistical testing
It is 1/169 = 0.005917, approx.
What events occur in a cell going through telophase?
A nuclear membrane forms around the two daughter nuclei, the chromosomes go back to chromatin, the nucleous reforms, the mitotic spindle disappears, and the cytoplasm divides.
What is the probability that a die will land ona certain number?
There are six sides to a die. If we assume that all sides have the same chance of coming up then each side has a probability of 1 in 6 or 1/6. This is the theoretical probability.
Of course if the die is not perfectly shaped and weighted then it will not be 'fair' and the sides will have different probabilities of coming up.
What factor can a researcher control that will reduce the standard error?
In any experimental test there are factors that can contribute to error. For example in a biochemical test if you add the compounds with a faulty pippete then the amount of reagent per sample would vary and that would contribute to a noisy measurement. So by controlling experimental conditions as best as one is able error can be reduced. Never completely eliminated, but managed.
The probability of event A is the number of ways event A can occur divided by the total number of possible outcomes. For example, the number of ways you can role a single die is 6, the number of ways to get an even number (2,4, or 6) is 3. So the probability of an even number is 3/6 or .5
Out of 28295 people 21786 are male what is the probability they will be chosen?
The answer depends on how many are chosen at a time.
Usually applied to accounting, so I will answer based on accounting. If a company has 3000 accounts numbered from 1 to 3000. Lets say you are to look at 100 files. 3000/100 = 30, random sampling is not choosing every 30th account. That does not give an equal chance of selecting each file. File 101 would never be selected. This is a biased sampling. You must use some kind of random number generator to select the 100 files to look at so that every file has an equal probability of being reviewed.
How can probability be used to make predictions?
Well... with what I learned from Mrs. Franks, mt math teacher, she said for weather. For example there with be a probability of 75 degrees today.
What is the probability of NOT getting 1 or 2 on a 6 sided dice when rolled 15 times?
P(1or2)=2/6=1/3
P(4<=x<=6)=2/3
when rolled 15 times
(2/3)^15=.0023